Suppose you have a calculator that has only the operation of division along with the ten numerals, an equal sign, and memory.
Explain how you could find 0.5% of 5600. (note: there are no decimal or multiplication operators)

Answer by khwang(438)

On way:
Step 1) Get rid of %, note % means 1/100,
so 0.5% of 5600 = 0.5 times 56 (cancel two zeros of 5600)
Step 2) Note 0.5 = 5/10 = 1/2 , ie 0.5 means one half.
Step 3) Key in 56 and then it divide it by 2 , you will get the answer.
That's the whole simple story. (No decimal & no multiplication operators)
Kenny